two automobiles travel at the rate of 50 mph and 30 mph respectively. The faster travels 2 hours more than the other and goes 140 miles farther. Find the distance traveled by each.

Answer by richwmiller(17219)

50*(t+2)=d+140
30t=d
50*(t+2)=30t+140
50t+100=30t+140
20t=40
t=2
50*4=200 miles at 50 mph for 4 hours
30*2=60 miles at 30 mph for 2 hours